OpenStudy (anonymous):

The volum eof an object is 8.12X10^-3 m^3. Calculate the volume in cubic feet. (2.54cm=1inch, 12.0 inches=1 foot)

OpenStudy (anonymous):

@zepdrix @DanJS

OpenStudy (danjs):

you just need to convert m^3 to ft^3 ..

OpenStudy (anonymous):

but how do i get m^3 removed.

OpenStudy (danjs):

i guess you have to use those easy to remember ratios they gave in parenthesis to convert between SI and US units, no looking up meters to feet directly

OpenStudy (danjs):

this way, not to difficult, you still want cubed units, a volume

OpenStudy (anonymous):

8.12X10^-3 m^3 (... i do not understadn this but I know I need a ratio that will cancle M6

OpenStudy (anonymous):

m^3*

OpenStudy (danjs):



OpenStudy (danjs):

start with the given, and multiply by ratios of 1 to change units cubed units to linear units is just like that, you have to distribute all those ^3 into the number and the unit

OpenStudy (anonymous):

how di du got the first ratio? that's usually the hard part for me

OpenStudy (anonymous):

1cm^3/ 10^-2 m^3

OpenStudy (danjs):

10^-2 m = 1 cm to change to cubed units, cube everything, constants and units [10^-2 m] ^3 = [1 cm]^3 10^-6 m^3 = 1^3 cm^3

OpenStudy (danjs):

see that

OpenStudy (anonymous):

okay I understand it now! Ty hope I will able to do these by myself. thnks

OpenStudy (danjs):

yeah just think of a conversion ratio as an equation 2.54 cm = 1 in to change to another dimension, just manipulate the equation with powers whole numbers to go up, and fractions to go down dimension

OpenStudy (danjs):

2.54^3 cm^3 = 1^3 in^3

OpenStudy (anonymous):

okay so anythin gthat would help cancel out the ratio given. try to mold the next ratio to do so. Alright thanks a bunch

OpenStudy (danjs):

to go from length to area, you use power 2, exe...

OpenStudy (danjs):

i always just start with the given thing then notice how units cancel, which tells you what way to put the fraction, m^3 has to be in the bottom to cancel out. work towards the unit you need

OpenStudy (anonymous):

alright I will practice more problems!

OpenStudy (danjs):